What is the angle of refraction in a medium if the angle of incidence in air (n=1.00) is 48˚ and the index of refraction of the medium is 1.58?

1
Expert's answer
2021-05-30T13:31:31-0400

"\frac{\sin48°}{\sin\ x°}=\frac{n_2}{n_1}\to x=\sin^{-1}(\frac{\sin48°}{n_2})=\sin^{-1}(\frac{\sin48°}{1.58})= 28°"